NOvA: mostly pretty smooth running this week. Experiment planning work on Far Detector and electronics. Getting ready for the power outage on 18th. Worked with ISD to schedule MINOS chiller cleaning on the 18th, same day as the power outage, to minimize effect of any potential cooling issues on detector. ArgonCube: moved detector electronics racks from LArTF to MSB and Underground. Thanks to all the people involved, especially Tom Wicks’ crew. MAGIS: laser room work in the MINOS backroom continues. Still talking about how and when to schedule power being down to the backroom racks so we can relocate the pump controllers. Also, work installing a second access door so tornado shelter can be reached once the laser room is complete. SBC: installed a roof over area planned for experiment.
Elevator maintenance should be on 29th, end of month.